OpenStudy (therealplatypus):

A recipe that makes 9 servings calls for 2 1/4 cups of juice. How many cups of juice are there per serving?

Vocaloid (vocaloid):

find: total number of cups/number of servings

OpenStudy (therealplatypus):

How?

Vocaloid (vocaloid):

read the problem

OpenStudy (therealplatypus):

yeah I did

Vocaloid (vocaloid):

good, then do the division

OpenStudy (therealplatypus):

2 1/4 divided by 9?

Vocaloid (vocaloid):

yeah

OpenStudy (therealplatypus):

k

OpenStudy (therealplatypus):

2.25?

OpenStudy (therealplatypus):

2 and one quarter is what I got

OpenStudy (therealplatypus):

woops I mean 4

OpenStudy (therealplatypus):

HELP

OpenStudy (therealplatypus):

VOCALOID I WILL MEDAL AND FAN JUST HELP

OpenStudy (therealplatypus):

@Vocaloid

OpenStudy (therealplatypus):

@Vocaloid

Vocaloid (vocaloid):

just convert 2 and 1/4 to a decimal 2.25/9 = ?

OpenStudy (therealplatypus):

I did that and I got .25

OpenStudy (therealplatypus):

but the choices it gives me are 3/8 , 2/3 , 1 1/3 , and 3

OpenStudy (therealplatypus):

@Vocaloid what am i doing wrong?

Vocaloid (vocaloid):

hm, weird

OpenStudy (therealplatypus):

oh wait

OpenStudy (therealplatypus):

I was looking at the wrong question

OpenStudy (therealplatypus):

so the answer is .25 (1/4) ?

OpenStudy (therealplatypus):

thats an option

Vocaloid (vocaloid):

yeah

OpenStudy (therealplatypus):

k I have 5 more like this

OpenStudy (therealplatypus):

A recipe that makes 6 servings calls for 1/2 cup of butter. Nancy modifies the recipe so that it can serve 8. How many cups of butter does Nancy need?

OpenStudy (therealplatypus):

@Vocaloid

OpenStudy (therealplatypus):

ugh I give up

OpenStudy (anonymous):

\[\frac{1}{9} \left(2+\frac{1}{4}\right)=\frac{1}{4} \]